Home page News page : Last Updated 27-Nov-2007 Project Description page Data page Sensors page Contacts page Calendar page Join page Links to related sites Partner Area
 

In April 2008 the ANIMATE network Observatories became part of the EuroSITES European Ocean Observatory Network.
Data shown here and subsequent data are available through the EuroSITES website www.eurosites.info
The ANIMATE site will not be updated further.

 

15th October 2007

CIS: The telemetry buoy came adrift shortly after 1500 utc of the 14th October and although it is still sending its position it is no longer sending data from the rest of the mooring.

 

9th July 2007 CIS: The mooring was successfully serviced by the research vessel Maria S Merian, and data from 12 microcats and the SAMI Carbon Dioxide Sendor are being regularly received via satellite.

 

22-June-2007 The redesigned PAP mooring was successfully deployed from the vessel Celtic Explorer

 

30-Aug-2006 Johannes Karstensen has sent this report:

I am happy to report that the CIS mooring was successfully re-deployed yesterday 29. Aug. 2006 and for the very first time we have access to on-line biogeochemical data (pCO2 and fluorescence) from the central Irminger Sea!!

According to the pressure data the mooring may be about 5 m too deep in the water (biogeochemical frame at 47m while target was 42m) - however, this could be the effect of the currents and we will know this only after a number of data points have been recorded.

The deployment went smoothly given the great support from the DISCOVERY personal and good weather conditions. To prevent the heavy corrosion that we experienced during CIS5 specific care was taken to 'seal' all places where different metals come into contact with Teflon foil or similar.

In particular this deployment is a success for Andreas Pinck and not to forget Michael Busack who worked hard to made the telemetry happen. Cross fingers that the CIS telemetry will be a follow up of the MOVE M1 which is online since about 16 month.


30-June-2006
The following message has been received from Dr Richard Lampitt

I am now at the PAP site on board Discovery and have to report the very bad news that all three of the upper ocean moorings have been destroyed by long line fishing. We have only recovered the lower portions of the moorings and even down to 2000m we found the remains of long lines with severe damage to the mooring line still remaining.

Until three years ago I had never come across a long line on the moorings and then it was only lightweight equipment and did not damage the mooring. The material we recovered yesterday was very much heavier and moorings in future will have to be much more robust to withstand them.

As a result of this and the liklihood it will happen again I have decided not to deploy the near surface morings planned for this cruise. I shall be discussing how they may be strengthened on my return to Southampton on 9th July and ways in which they may be deployed this year.

We have all lost from this experience both in terms of science, time and hardware and I am sorry not to be able to bring you a few items of optimism. I hope we will recover the deep sediment trap mooring tomorrow unless of course the fishermen are working as deep as 3000m.


19th October 2005

The drifting part of the 2005 PAP telemetry mooring has been recovered by the RRS Discovery which currently undertaking the annual Atlantic Meridional Transect cruise (AMT). Many thanks to the scientists and crew.


2nd October 2005

Short report of servicing of CIS Mooring by RV Pelagia by Michael Busack

The interdisciplinary CIS (Central Irminger Sea) mooring is funded by the European project MERSEA. The recovery of the 2745m long mooring took place on 16 September 2005. It included :

- 14 SBE37 (MicroCats), salinity and temperature recorders, 6 of them with pressure sensors, in several depths down to 1500m

- 2 ADCPs at 151m (Workhorse 300kHz up-looking, LongRanger 75kHz down-looking)

- a sensor frame at 40m with a nutrient sensor, fluroescense, pCO2 and also a MicroCat for T,S and P

- two RCM-8 current meters and a 21cup sediment trap

- and a small telemetry bouy at the surface for realtime data transmissions (this failed shortly after deployment on the Charles Darwin 161 in September 2004).

All instruments worked well and recorded data.

For the new deployment two days later it was necessary to check, service and calibrate 6 of the 14 MicroCats, 8 new ones came from Kiel. The two ADCPs got new Batteries and were also re-deloyed in the new mooring. For the sensor frame at 40m new nutrient, flurenscense and pCO2 sensors had been provided by the project partners. A sediment trap is not re-deployed, two new RCM-8 arrived also from Kiel.

The new mooring was deployed on the 19 September again with a telemetry buoy, which now includes an additional temperature sensor at the surface. Otherwise the mooring is identical to the recovered one. First data received indicated functioning of the telemetry system but only to 40m depth.


July 2005
RRS Discovery has successfully serviced the PAP mooring.
March 2005
RV Poseidon has successfully serviced the ESTOC mooring.
9-17th December 2004
RV Poseidon has successfully serviced the ESTOC mooring.
6th December 2004
Johannes Karstensen has issued a report on the methodology for calibrating and correcting microcat, ADCP and RCM data.
The report can be viewed by clicking HERE.
17th November 2004
The telemetry buoy of the ESTOC mooring has become adrift.
Fortunately the RV Poseidon will shortly be passing the site, prior to a cruise to service ESTOC so the drifting parts should be soon recovered and the reason for the mooring failure acertained.
23rd September 2004
Work at CIS during CD 161:
After the 24-hour delay due to bad weather at mooring K81 in the LabSea, we caught up 12 h to reach CIS on 17.09.2004 in the afternoon; however, heavy swell from the NE and increasing winds from SW prevented recovery of CIS on the same day. As the forecast would allow recovery next day we decided for that option, and for the time being obtained the 300 m CTD calibration cast with the new WetLab fluorometer attached to the rosette, and samples for O2, CO2, nutrients, and filtration taken with high vertical resolution around the moored sensor frame’s nominal depth of 40 m.
As forecasted for next morning, 18.09.2004, both swell and wind had decreased which made it possible to recover mooring V434-03/CIS-03 without problems after its deployment earlier this year in May from the Icelandic R/V B SAEMUNDSSON. Downloading and inspection of recorded data showed that all physical instruments, microcats, WH ADCP, LR ADPC and RCM8 had recorded reasonable data, and that the trap had the expected 5 cups closed. The CO2 sensor SAMI #25 was the only chemical sensor that had been implemented for that 4-months period. It has recorded reasonable data for 3 weeks, but has failed thereafter. Opening the instrument and investigation showed that one of the thin pipes which allows for CO2 to diffuse into its indicator solution, had come loose from its fixing point. It is unclear how this could happen, but this also led to a loss of chemicals. As no spare instrument had been delivered to St. Johns prior to the cruise, SAMI #25 was repaired, again inspected to work correctly and prepaired for re-deployment, however increasing the recording interval from 2 h to 3 h to partly compensate for the loss of chemicals and saving battery capacity.
According to the the forecast for the next two days, wind should increase up to gale 8 and swell again become heavy. In view of the tight schedule, this forced us to concentrate preparations for redeployment to the minimum: spool wires; maintain the physical sensors to be redeployed; replace batteries in the new WetLab sensor and reset its sampling rate after the calibration cast; check the new telemetry with the microcats and with the CO2 sensor. The new McLane trap and the new nutrient sensor and two acoustic releasers were already ready to launch. Also, a CTD cast with sampling for O2, CO2, and nutrients down to the bottom for comparison was taken. However, due to the tight schedule, no microcats could be attached for in-situ calibration. Preparations were finished late evening, and deployment started after rest of 6 h early morning of 19.09.2004. During deployment, final checks of the telemetry were performed and showed no failures. After having slipped the anchor, the telemetry was spotted to stay fixed on position by 10:15. We left CIS and while speeding up, launched an APEX float with oxygen sensor. As the final diagnosis of the PROVOR float to be launched still was negative, it was not deployed, but will be returned to the manufacturer for repair under warranty.

13th September 2004
RV Charles Darwin has been repaired and after a successful sea trial resumed cruise 161/162 on Saturday 11th September. Work has been rescheduled and the ship is estimating arrival at Reykjavik in the morning of 22nd September.


3rd September 2004
RV Charles Darwin set out from St Johns, Newfoundland, on the 28th August to service the CIS moorings but has developed engine trouble and is returning to St Johns.
It may not be possible for all the work to be completed on this cruise so alternatives are being explored.


Click here to see a beautiful image of dust blowing from Africa across the ESTOC site (26-Aug-2004).


23rd June 2004
RV Charles Darwin is currently on site and has redeployed the PAP telemetry mooring.


17th May 2004
RV BJARNI SAEMUNDSSON has successfully serviced the CIS mooring.

The folowing short report was issued by Thomas Müller:
The B SAEMUNDSSON with PS Hedinn Valdimarsson and two groups from MRI and IFM-GEOMAR successfully has recovered both, CIS1 and CIS2, and launched a single mooring that for the first time combines both, the telemetry and all other sensors.
The weather conditions on Friday, 15th May, initially were not good, with up to 8 m waves and high swell, but improved over the time until Sunday, 17th May. Thus, one technician from MRI was hurt due to a breaking lashing, fortunately not too severely. One of the recovered mooring wires broke while at the the surface, and this part needed to be found and picked up again. Finally, however, all instruments were on deck, and in rather good condition.

>From a short German written report the data return and re-deployment info:
-MCs: data downloaded, look ok; re-deployed
-2 ADCPs: data downloaded, look ok; re-deployed
-2 RCM8: both have the same amount of data on the data storage unit DSU; to be downloaded in Kiel; instruments need re-furbishment; 2 back-up RCM8 deployed
-NASE2, nutrient sensor: data downloaded; quality control remains to ICCM; no re-deployment because of lacking back-up sensor; sensor to be returned to ICCM
-SAMI, CO2: data presently to be downloaded; re-placed by back-up sensor
-HS2, Fluorescence: shutter broken off; sensor worked on deck; no data downloaded because of software problems; no re-deployment because of broken shutter; sensor to be returned to SOC; no re-placement because of lacking back-up sensor
-trap: no information yet; filled cups should have been removed; trap should have been re-deployed with remaining cups


28th April 2004
RV POSEIDON has successfully serviced the ESTOC mooring but unfortunateley only the top 5 microcats are currently sending real-time data.


30th March 2004
ANIMATE 2004 1st Quarter Report
The first quarter is very much the time for planning. In January the project partners attended a meeting in Southampton where results gathered so far were presented and discussed and the experiences over the previous two years used to plan the deployments for the coming year. But even though it is the winter season there are still on water activities. In February Michael Busack managed to change the telemetry buoy at the ESTOC mooring and restore satellite communication and in March a Danish trawler support tanker, EMMA, recovered the parts of the CIS2 mooring which had been adrift since mid November, so a lot has already been achieved and the first 2004 ANIMATE cruise is still a fortnight away! And the latest news is that the ANIMATE 2nd annual report has been finalised and submitted and is now available on the partner area. The annual report summarises the achievements so far, but also shows how much more can be achieved before the end of the project in November 2004. It has been a good start to the year, but there is still much to do.


24th March 2004
Part of the CIS telemetry mooring broke free in November and we now have news that a Danish trawler supply vessel has recovered the telemetry mooring and 5 microcats.


13th February 2004
Michael Busack from Ifm-GEOMAR has been able to change the telemetry bouy on the ESTOC mooring so that we are again receiving data from the Microcats on sea temperature and salinity down to 945m.


27th January 2004
Technical Meeting 27th Janurary
Partner Meeting 28/29 January 2004 at Southampton Oceanography Centre.


5th January 2004
Data received from ESTOC during the last week of December has failed quality checks and since midnight 4th January no data at all has beeng delivered.


End of Year Roundup

Well what a roller-coaster year this has been with ups and downs ..... but (fingers crossed) it is ending on some very positive ups!

January - It was decided to redesign the buoyancy associated with the sensor frame to make it easier to deploy, and to design and make conductive swivels so that the sensor package mooring could be incorporated with the telemetry link mooring.

February - The CIS telemetry mooring went adrift in early February and then worse still the CIS sensor frame surfaced. When a ship in the region located the buoy which should have been sending out the sensor frame position it was no longer attached to the sensor frame, so all sensors have been lost. Johannes Karstensen has been able to use the data already collected at CIS to produce some interesting heat budget calculations, comparing the ANIMATE data with the General Ocean Turbulence Model using NCEP/NCAR reanalysis surface flux data. Graphs were placed on the CIS area of the website

March - The part of the PAP mooring which had been drifting since early December was finally recovered on the 1st March 2003 but because of the battering it had received during the 14 weeks it was adrift it revealed very little about why the mooring had failed in the first place.

April - ESTOC was redeployed on 12th April and the real-time data on temperature and salinity were soon available on the web site. Weather data is sent in real-time from DOLAN, which is only a few kilometres from ESTOC, via the ORBCOMM system. Air temperature, air pressure and relative humidity were all made available on the web site.

June - The SAMI CO2 sensor and a microcat were deployed at the DOLAN site . Data was sent in real-time from DOLAN but unfortunately the SAMI sensor was not functioning correctly but sea temperature and salinity at 10m depth were made available on the web site.

July - The PAP site was serviced. None of the failed microcat mooring was recovered but the sensor package was recovered with all the sensors together with the separate sediment trap mooring. Although the new telemetry buoy containing the satellite link was working at the start of the cruise it failed later tests so the usual 2 moorings for microcats and sensor package were combined into a single mooring and all sensors set to collect data in delayed mode only.

August - The CIS site was serviced. And although only the sediment traps were recovered from the previous deployment , a fully equipped sensor package mooring, satellite mooring and sediment traps were deployed. An additional air pressure sensor had been attached to the telemetry mooring but this was completely submerged within a few hours by a strong current effect and this seems to have caused both the air pressure sensor to fail and then the entire satellite telemetry link. The data will continue to be logged by the sensors and will be available as delayed-mode data.

September - ESTOC was serviced. A new SAMI CO2 sensor was deployed at the DOLAN site and the first CO2 data was received and plotted real-time on the ESTOC area of the website. The only change to the mooring design between deployments was to move the 30m ESTOC microcat sensor to 10m so the graphs and data base table which have been being populated since 12th April continue to be used.

October - An extra cruise was arranged to service the PAP mooring. During the season it has become apparent that the splice eyes whicha are used to terminate the steel cables are prone to corrosion. The ESTOC an CIS moorings have been redeployed with a different type of terminator and it is believed that the PAP mooring will be likely to fail unless these elements in the mooring are changed.

November - PAP was successfully serviced. Despite the lateness in the year the R.V. Poseidon science team and crew found enough of a weather window to recover and redeploy the PAP moorings and we are now receiving the temperature and salinity via satellite again. Every piece of good news seems to be tempered with some bad: it appears that the ARGOS watchdog for the CIS sensor package, and hence the sensor package it was attached to, are now adrift.

December - Now that the cruises are over for 2003, and before the detailed planning begins in earnest, there is a small respite. Some of the delayed mode data collected during the cruises has already been plotted as graphs which have been placed on the web site. The data will be quality checked and processed to be in user friendly formats and placed on the web site as soon as possible. The Data Manager is hoping to be overwhelmed by the data which will be exchanged at the next ANIMATE meeting, which will be in Southampton on 28th and 29th January. The considerable achievements so far of ANIMATE will be assessed and plans for capitalising on the lessons learned from both successes and setbacks will be discussed and applied to the 2004 ANIMATE program.

Merry Christmas
Fröhliches Weihnachten
Feliz Navidad
Glešileg jól
Joyeux Noėl



19/11/2003 -

Congratulations to the science team and crew of the research vessel Poseidon!

The mooring deployed at the PAP site during July has been recovered, the delayed mode data downloaded and the sensors serviced, recalibrated and a complete set of sensors redeployed. Poseidon is now returning to Bremerhaven and is expected in port on the 24th November.


03/11/2003 -
ESTOC mooring has been successfully serviced. Data are again being collected real-time via satellite.


28/08/2003 -
CIS redeployment The redeployment of the CIS moorings has been successfully completed by scientists and crew on the RV Poseidon.


08/07/2003 -
PAP redeployment is completed! The Microcats remaining in the broken PAP2 mooring did not surface when released and the satellite transmitter due to be deployed has failed so all the sensors and the available Microcats have been deployed but no real-time data will be available. The delayed mode data recovered will be processed and placed on the web site as soon as possible

Cruise Schedule

Click here for cruise schedule

Old News

Click here for all the older news stories

Project started: December 1st 2001

      Back to Top of page